a) Consider a merger among firms producing two complementary goods, computer (x) and memory card (Y). The consumers treat bota) Suppose that the X producer and the Y producer are independent. Solve for the Nash-Bertrand equilibrium in prices. Calculaa) Suppose that the X producer and the Y producer are independent. Solve for the Nash-Bertrand equilibrium in prices. Calculaa) Suppose that the X producer and the Y producer are independent. Solve for the Nash-Bertrand equilibrium in prices. Calculab) Now suppose that firm X and Y merge under a single ownership. Calculate the monopoly equilibrium prices, the quantity prodHi, can anyone solve the following questions a) and b) circled in red in a easier and simplied way ?. I have also posted the solution for this. Thanks
Answer #1

classmate Date Page bofit - Revenue - costs x= PxQ - 0 Px (X-Ps) (production is coolless) (Q = x gwen) [Ps - Pxtaly Px (4-Pn-classmate Nay, les- Pataly Date Page X 2 3 Ps= 22 Q = X-Ps (given) Q = *-28 - z Calculating Profits for seller of a & y 2 Te